OpenStudy (kiara_ahquin12):

(1 pt) Points Y and Z are the same distance from point X on a number line. The coordinate of point X is –7 and the coordinate of point Y is –1. What is the coordinate of point Z? (1 pt) Points Y and Z are the same distance from point X on a number line. The coordinate of point X is –7 and the coordinate of point Y is –1. What is the coordinate of point Z?

OpenStudy (kiara_ahquin12):

plz help

OpenStudy (jdoe0001):

http://putnam.k12tn.net/cms/Probo/Coordplane.png <--- notice the cartesian grid and move that many units on the given direction, and you should be able to see where it will land you

OpenStudy (royalranger):



OpenStudy (kiara_ahquin12):

is it 5 units

OpenStudy (kiara_ahquin12):

-5

OpenStudy (royalranger):

-7 - (-1) =?

OpenStudy (kiara_ahquin12):

-6 im not quite sure

OpenStudy (kiara_ahquin12):

is it correct

OpenStudy (kiara_ahquin12):

so just 6 not -6

OpenStudy (royalranger):

I guess that just confused you, never mind it. what is: -7 + (-6) =?

OpenStudy (kiara_ahquin12):

1 i think

OpenStudy (royalranger):

No, what is a negative plus a negative?

OpenStudy (kiara_ahquin12):

negative

OpenStudy (kiara_ahquin12):

i have no idea can you just kinda give a the awnser lol

OpenStudy (royalranger):

Sure i could, but i cant give it to you on a quiz or test.

OpenStudy (kiara_ahquin12):

its hw

OpenStudy (royalranger):

Nevermind... Alright so like you said, a negative plus a negative = a negative. Now that we know the final answer will be negative. Lets just get rid of the negatives to make life easier: What is : 7 + 6 =?

OpenStudy (kiara_ahquin12):

13

OpenStudy (kiara_ahquin12):

wait is it -13

OpenStudy (royalranger):

Yep.

OpenStudy (kiara_ahquin12):

okay that was my first awnser i just wanted to make sure it was correct thank you so much

OpenStudy (royalranger):

No problem.